OpenStudy (anonymous):

The population of a local species of mosquitos can be found using an infinite geometric series where a1 = 740 and the common ratio is one sixth. Write the sum in sigma notation, and calculate the sum (if possible) that will be the upper limit of this population.

OpenStudy (anonymous):

\[\sum_{i=1 }^{ ∞}740(\frac{ 1 }{ 6})\] but is the power to 1/6 ^i or i =1 also is it divergent? or does it have a sum im confused on the rest

OpenStudy (anonymous):

@ganeshie8 @iambatman @cwrw238

ganeshie8 (ganeshie8):

are you asking if the exponent would be \(\large i\) or \(\large i-1\) ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

no i = 1 though not -

ganeshie8 (ganeshie8):

if \(i\) starts from \(1\), the first term must evaluate to 740 right ?

ganeshie8 (ganeshie8):

\[\large \sum_{i=1 }^{ \infty}740\left(\frac{ 1 }{ 6}\right)^{i-1} \]

ganeshie8 (ganeshie8):

the exponent has to be \(\large i-1\) otherwise you wont get the first term as 740

OpenStudy (anonymous):

oh yeah my bad sorry couldn't see the minus clearly.

OpenStudy (anonymous):

now how do I find if its divergent? thats what Im most confused on

ganeshie8 (ganeshie8):

Notice that the common ratio is 1/6 which is less than 1 so the geometric series converges

ganeshie8 (ganeshie8):

\[\large \text{if } |r| < 1, \text{ then the series converges} \]

ganeshie8 (ganeshie8):

use the infinite geometric series formula to evaluate the sum

ganeshie8 (ganeshie8):

\[\large \text{infinite sum} = \dfrac{a}{1-r}\]

OpenStudy (anonymous):

it is 888

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

888 is \(\large \color{red}{\checkmark}\) good job !!